# Industrial Internet of Things Intelligence Empowering Smart Manufacturing: A Literature Review

### 02 Dec 2023 v1

@@ VIII Conclusion
We have designed a hierarchical development architecture for IIoT intelligence to build a systematic understanding of how IIoT intelligence deals with the issues in the transformation process of smart manufacturing. The architecture consists of equipment layer, networking layer, software layer, modeling layer, and analysis and optimization layer. The equipment layer builds the foundation of IIoT intelligence. The networking layer realizes ubiquitous connection among human-machine-thing. The software layer provides digital expression of industrial processes. The modeling layer achieves accurately modeling and interoperability improvement. The analysis and optimization layer supports case-by-case algorithm design to analyze industrial data and optimize industrial processes. In addition, we also closely observe the lighthouse factory to demonstrate the positive effects of technologies at each layer.
@@
Following that, we identify seven kinds of technologies that accelerate the development of IIoT intelligence, including industrial robots, machine vision, networking technologies, digital twins, deep learning, smart hardware, cloud/edge computing. Industrial robots can become intelligent labors to overtake dirty and dangerous tasks. Machine vision has potential to be the eyes of industry. Networking facilitates ubiquitous connectivity. Digital twins connect cyberspace and physical spaces. Smart hardware builds the foundation of intelligence. Cloud/Edge computing inspires novel business and deployment.

### 22 Feb 2024 v2

@@ IX Conclusion
We propose a hierarchical development architecture for IIoT intelligence to build a systematic understanding of how IIoT intelligence deals with the issues in the transformation process of smart manufacturing. The architecture consists of equipment layer, networking layer, software layer, modeling layer, and analysis and optimization layer. The equipment layer builds the foundation of IIoT intelligence. The networking layer realizes ubiquitous connection among human-machine-thing. The software layer provides digital expression of industrial processes. The modeling layer achieves accurately modeling and interoperability improvement. The analysis and optimization layer supports case-by-case algorithm design to analyze industrial data and optimize industrial processes. In addition, we also closely observe the lighthouse factory to demonstrate the positive effects of technologies at each layer.
@@
Following that, we identify seven kinds of technologies that accelerate the development of IIoT intelligence, including industrial robots, machine vision, networking technologies, digital twins, deep learning, smart hardware, cloud/edge computing. Subsequently, we analyzes the ethical implications and environmental impacts of adopting IIoT intelligence in manufacturing.